# SuperARC: A Test for General and Super Intelligence Based on First Principles of Recursion Theory and Algorithmic Probability

### v1

@@ 4.1.3 Compression as Comprehension and Prediction
It is important to clarify possible misinterpretation of the meaning of the word “compression” as used in our framework. In machine learning and cognitive science, feature selection involves identifying the most relevant variables or attributes that contribute to predictive modelling. This process reduces dimensionality, focusing on the most informative aspects of data. It is, of course, a compression approach, but just a part of the one we intend to refer to. Model compression in our framework also refers to simplifying a model without significantly compromising its performance. It involves reducing the complexity of the model, often leading to better generalisation and greater efficiency. It is, therefore, related to model building and data pre-processing (automatically done by the model).
@@ 4.1.4 An updated definition of Intelligence
Based on the AIT principles described, we can update our definition of intelligence to read as follows :
Intelligence is the ability to create a computable model that effectively (as losslessly as possible) explains any given data, where greater intelligence corresponds to models with lower algorithmic complexities.
Using algorithmic complexity as a measure of model compactness provides an agnostic and absolute metric, as its value corresponds to the shortest possible program capable of reproducing a given dataset. This establishes a universal definition of intelligence, serving as both a theoretical and practical upper bound for its highest possible level.
Unlike standard tests that assess intelligence based on predefined ‘correct’ answers–inevitably influenced by subjective notions of correctness–we shift the focus to identifying the shortest possible explanation for a given dataset. In our framework, correctness is defined purely as the ability to reproduce the data exactly (losslessly), while intelligence is measured by achieving this with the most concise program or formula.
As a result, the SuperARC framework accommodates any type of data as input-output pairs, requiring only that a complexity-based metric be predefined. To achieve this, in preliminary explorations we will approximate algorithmic complexity by methods like LZW and ZIP which are more closely related to Shannon Entropy [34], but we will also use the Block Decomposition Method (BDM) as our gold-standard approach. The latter is based upon the Coding Theorem Method (CTM) – a direct consequence of Algorithmic Probability – and therefore also able to quantify algorithmic probability.
In other words, we provide a theoretical underpinning which suggests that an intelligent agent must excel at both compression (abstraction) and prediction (planning) to be considered truly intelligent. Designing tests that measure these abilities can lead to a more nuanced and computationally grounded understanding of intelligence, applicable to both human cognition and artificial intelligence systems.
@@ 4.2 A Neurosymbolic Approach to a Superintelligence Benchmark
Using the principles of classical information theory, the Block Decomposition Method (BDM) or BDM combines the calculation of the global Shannon Entropy rate of the object with local estimations to algorithmic complexity of smaller blocks into which the object is decomposed for which values are found in a pre-computed database of direct approximations of algorithmic probability. One way to think of BDM is by depicting it as a Deep Learning Transformer which aims to build a predictor that maximises the probability of being correct in explaining the data by looking for long-range and short-range correlations. The difference, in this case, is that long-range correlations are covered by Shannon Entropy (not fundamentally different from Transformers) but short-term correlations are estimated using the principles of algorithmic probability through the Coding Theorem [62, 63, 64, 59]. This therefore combines the two best methods for statistical and algorithmic inference.
@@
Formally, let $x$ be a string divided into blocks $x_{i}$, with $x=x_{1}\oplus x_{2}\oplus\dots\oplus x_{n}$, where $\oplus$ denotes a concatenation operator. The BDM complexity of $x$, denoted $\text{BDM}(x)$, is:
$\text{BDM}(x)=\sum_{i=1}^{n}\text{CTM}(x_{i})+\log n$
@@
  • $\log n$ is a correction factor accounting for the interactions between the blocks.
The Coding Theorem Method (CTM) is a method based on the Coding Theorem and Algorithmic Probability [26, 58], which connects classical probability to algorithmic complexity [63, 64, 62]. The CTM maps sets of micro programs (e.g., small Turing machines) to small assembly objects for which it can empirically estimate the algorithmic probability of an object, such as a time series, based on the following relationship [68]:
$K(s)\approx-\log P(s)$
@@
  • $K(s)$ is the algorithmic complexity of string $s$.
  • $P(s)$ is the algorithmic probability of string $s$, as defined by the universal distribution.
CTM produces and stores the set of Gödel numbers that correspond to all the programs that compute an object, such as an integer sequence, up to the given digit or any other recursively describable [63, 64]. Each program can then be uncompressed from its unique (Gödel) number and run to produce the next digit for predictive purposes with the programs themselves the abstract future-planning models. While CTM operates by brute force, BDM then leverages the pre-computed distributions from those small computer programs to stitch them together according to the rules of information theory to guide the search of the best sequence of programs explaining larger objects.
On the one hand, CTM provides an approximation to algorithmic probability $P(s)$ by connecting the empirical frequency of occurrence of an object produced by a random computer programme with its algorithmic complexity $K(s)$, using the relation $K(s)\approx-\log P(s)$ and also keeps track of the set of programs that generated the original object, hence identifying the mechanistic generators. On the other hand, BDM offers a method to map the micro programs produced by CTM to their corresponding pieces from the larger object to explain by decomposing the original object into smaller blocks for which micro programs have been found by CTM with a correction factor for block interactions (e.g. repetitions).
In the context of integer sequences, the BDM and CTM can be applied to test both:
@@
  • Prediction as planning: Using the BDM complexity as a proxy for the time series’ regularity, one can infer the predictability of future values. Lower BDM complexity implies a simpler underlying structure, which can help in forecasting future elements of the series–which is similar to how algorithmic probability and Levin’s universal distribution can be used for predictive modelling. This is related to planning, because once several program pathways are identified, one can verify each against the next token and update the program set (by discarding those programs that did not fit the next token) while keeping the shortest program criterion.
@@ 4.2.1 Why CTM and BDM as standard for abstraction and planning
  • CTM helps identify the set of candidate underlying generative mechanisms and provide a set of models from which it can actively predict future values by running it further into the future providing a set of projections. CTM forecasting requires an iterative refinement process in which multiple possible generative programmes are tested and updated. CTM can help select the most likely program candidates from CTM by favouring those with lower complexity in accordance with the principles of algorithmic probability.
@@ 4.3 Comprehension via Algorithmic Probability
19:$\mathcal{T}_{positive}\Leftarrow\alpha(\mathcal{T}-\min(\mathcal{T}))+\epsilon$triangleright Affine transformation to ensure all values are positive and differences remain proportional
@@
It should also be noticed that CTM/BDM is not purely a brute-force approach, requires no previous data and its current implementation required orders of magnitude less computational power. While CTM alone would be a brute-force approach that looks for the shortest computer programs explaining the data, BDM combines it with traditional pattern-matching, meaning that CTM/BDM meets the best of both world in the middle, right at the fine balance between what traditional Machine Learning and Deep Learning approaches are able to achieve while combining it with optimal Bayesian causal inference [34, 69]. We have called this approach Algorithmic Information Dynamics [71, 58, 59].
@@ 5.1 Next-digit Prediction Task with Binary and Non-binary Sequences
Figure 3 shows an increase in complexity as was expected given the design of each group of generated sequences. The plot suggests that BDM can capture (and can generate) better complexity and randomness, since its values increase more consistently as complexity increases, unlike other measures. Shannon-entropy-based measures (and cognates) can account for statistical randomness only. Compression algorithms, for example, decrease as complexity increases, becoming more difficult to find regularities and increasing compression length as a function of complexity growth.
@@ 5.3 Code Generation Task with Non-binary Sequences
According to the results ( 8 top), the vast majority of correct cases are print failing to compress the sequences. This indicates that in most instances where the system correctly identifies a sequence, it does so by simply outputting the sequence as is, without any attempt at compression.
@@
Figure 9 (bottom) shows how no-compression generally increases with complexity, except for Mathematica, where the no-compression percentage is lower at complexity level 2 than at level 1. This happened because Mathematica has the capacity to computationally replicate several well-studied and known sequences of numbers. This capacity leads to shorter code at complexity level 2. However, at complexity level 3 the trend aligns with other languages, showing direct proportionality between complexity and no-compression.
@@ 6 SuperARC-seq
Based on the previous experiments, it is possible to characterise one test directly related to the SuperARC framework: the SuperARC-seq. The objective of this test is to quantify intelligence and related cognitive capacities, specifically, reasoning and comprehension, drawing inspiration from the work in [74] and the theoretical and empirical studies here introduced. As mentioned, this test is grounded in one of the fundamental cognitive tasks: recognising patterns and evaluating the complexity of finite sequences, which inherently requires a level of understanding in order to provide a meaningful explanation. In our experiment, we generated short binary sequences and tasked several advanced LLMs with deriving a formula for each sequence capable of reproducing the target sequence.
@@
    Ordinal: The model provided a mapping based on the indices where “1”s occur in the sequence. This response reflects an attempt by the model to analyse and map some logical structure to the sequence, making it more valuable than simply reproducing it verbatim.
@@
For any given tested model, the percentages of results belonging to each group can be combined as a vector of results, $\rho=[\%_{c,np,no},\%_{c,o},\%_{c,p},\%_{i}]$, such that $\sum\rho_{i}=1$ as the percentages will be represented in the range [0,1] to resemble probabilities. We know, beforehand, that the best performing model would be one with $\rho_{best}=[1,0,0,0]$. Thus, a first possible test would be to check the difference between correct and incorrect percentages.
$\varphi_{a}=\sum_{i=1}^{3}\rho_{i}-\rho_{4},$ (1)
which would range from -1 to 1 for models that are not able to reproduce any sequence to models which perfectly reconstruct the sequences, respectively. However, this only accounts for the ability of LLMs to reproduce the initial sequence (planning) but not for their compression capabilities. To account for the latter, let us assume that the best possible algorithm for each element of the data set is $\mathcal{B}_{k,j}$, such that $\mathcal{B}_{k,j}()=D_{k,encoded}[j]$, and here the algorithm does not have a particular input, similar to the definition of algorithmic complexity. Thus:
@@
due to the information non-increase theorem and to the fact that no inputs were used in the function. In addition, it is clear that $K(\mathcal{B}_{k,j})\leq K(R_{k,j})$ since $\mathcal{B}_{k,j}$ is the shortest possible programme to reproduce $D_{k,encoded}[j]$. Thus, from equation 2:
$K(D_{k,encoded}[j])\leq K(R_{k,j})$ (3)
This suggests that the ratio $K(D_{k,encoded}[j])/K(R_{k,j})\leq 1$ could be used as a weight to indicate how well the model achieves compression and comprehension, with the best possible value equal to 1 representing a perfect ratio scenario. Since we have several algorithms classified under each of the four types (according to their structure), instead of using the individual ratios, we shall use the harmonic mean per type, defined as:
$\delta_{k}=\frac{n_{k}}{\displaystyle\sum_{j=1}^{n_{k}}\frac{K(R_{k,j})}{K(D_{k,encoded}[j])}}\text{ for }R_{k,j}\text{ of type }k,$ (4)
@@
$\varphi_{b}=\sum_{i=1}^{3}\delta_{i}\rho_{i}-\rho_{4}.$ (5)
Deliberately, instead of calculating the harmonic average of complexity ratios for formulas of incorrect type ($\delta_{4}$), we penalise the test the most and set the weight $\delta_{4}$ to -1. Besides, we want to privilege models that do not simply copy or provide ordinal mappings of the input sequences. Thus we can attribute higher weights to types that are correct and do not copy nor print the results. We also want to give more weight to programs that provide ordinal mappings when compared to print cases. Then, considering a power-law weighting strategy, the final test metric is:
$\varphi=\delta_{1}\rho_{1}+\frac{\delta_{2}\rho_{2}}{2}+\frac{\delta_{3}\rho_{3}}{4}-\rho_{4}.$ (6)
It can be seen that $\varphi\in[-1,1]$ encompasses different behaviours. For example, $\varphi\in[0,0.25]$ if only print-type models are outputted. Also, $\varphi\in[0,0.5]$ if only ordinal-like formulas are created. Finally, $\varphi\in[0,1]$ in cases where the LLMs create formulas that are always correct, do not copy nor create ordinal mappings. The ranges will be populated with varying compression levels corresponding to the algorithms obtained. The test performance results for each model are calculated using equation 6 for $\mathcal{T}$ in Algorithm 1.
@@
As described, equation 6 tests for two features, compression via non-print computer programs and non-ordinal mathematical formulas to the input sequence, and prediction, by running all programs and all formulas to match each sequence digit, and penalising them when they did not represent an actual compressed model that generated a possible new digit of the sequence when run in reverse, i.e. when ‘decompressed’. The test formula assigns greater importance to correct cases that are not solutions of the type ‘print($s$)’ where $s$ is the sequence for which the AI system is asked for a model, given that a print model does not allow generalisation by prediction through simulation, as running a print command will only print up to the last digit. The same is true for what we call ‘ordinals‘, which is simply indicating the index of the non-zero non-one element in the binary sequence, meaning that, together with the ‘print’ case, the system failed in its attempts at abstracting features of the object. Finally, the formula punishes ordinal and print answers in a weighted fashion. The best performer can only reach a $\varphi$ of 1 while the lowest value is -1.

### v2

@@ 4.1.3 Compression as Comprehension and Prediction
It is important to clarify possible misinterpretation of the meaning of the word “compression” as used in our framework. In machine learning and cognitive science, feature selection involves identifying the most relevant variables or attributes that contribute to predictive modelling. This summarisation process reduces dimensionality, focusing on the most informative aspects of data. It is, of course, a compression approach, but just a part of the one we intend to refer to. Model compression in our framework also refers to simplifying a model without significantly compromising its performance. It involves reducing the complexity of the model, often leading to better generalisation and greater efficiency. It is, therefore, related to model building and data pre-processing (automatically done by the model).
@@ 4.1.4 An updated definition of Intelligence
Using algorithmic complexity as a measure of model compactness and optimal prediction provides an agnostic (human independent) quantitative metric, as its value corresponds to the shortest possible program capable of reproducing a given dataset and its optimal prediction value is governed by algorithmic probability. This can establish a universal definition of intelligence, serving as both a theoretical and a practical upper bound for the highest possible levels of compression such as model abstraction and prediction, which are believed to be fundamental features of intelligence.
Unlike standard tests that assess intelligence based on predefined ‘correct’ answers—inevitably influenced by subjective notions of correctness—we shift the focus to identifying the shortest possible explanation for a given dataset. In our framework, correctness is defined purely as the ability to reproduce the data exactly (losslessly), while intelligence is measured by achieving this with the most concise program or formula as a function of optimal prediction (via decompression).
As a result, the SuperARC framework accommodates any type of data as input-output pairs, requiring only that a complexity-based metric be predefined. To achieve this, we will approximate algorithmic complexity by methods like LZW and ZIP which are more closely related to Shannon Entropy [35], but we will also use the Block Decomposition Method (BDM) as our gold-standard approach that goes beyond statistical compression or statistical pattern-matching [61]. The latter is based upon the Coding Theorem Method (CTM)—a direct consequence of Algorithmic Probability [62].
In other words, we provide a theoretical underpinning which suggests that an intelligent agent must excel at both compression (abstraction) and prediction (planning) as a metric for (super) intelligence. Designing tests that measure these abilities can lead to a more nuanced and computationally grounded understanding of intelligence that is applicable to biological (e.g. animal), human cognition, and computational intelligence.
@@ 4.2 A Neurosymbolic Approach to a Superintelligence Benchmark
Using the principles of classical information theory, the Block Decomposition Method (BDM) combines the calculation of the global Shannon Entropy rate of the object with local estimations to algorithmic complexity of smaller blocks into which the object is decomposed for which values are found in a pre-computed database of direct approximations of algorithmic probability. One way to think of BDM is by depicting it as a Deep Learning Transformer which aims to build a predictor that maximises the probability of being correct in explaining the data by looking for long-range and short-range correlations. The difference, in this case, is that long-range correlations are covered by Shannon Entropy (not fundamentally different from Transformers) but short-term correlations are estimated using the principles of algorithmic probability through the Coding Theorem [63, 62, 64, 59]. This therefore combines the two best methods for statistical and algorithmic inference.
@@
Formally, let $x$ be a string divided into blocks $x_{i}$, with $x=x_{1}\oplus x_{2}\oplus\dots\oplus x_{n}$, where $\oplus$ denotes a concatenation operator. The BDM complexity of a string $x$, denoted by $\text{BDM}(x)$, is given by:
$\text{BDM}(x)=\sum_{i=1}^{n}\text{CTM}(x_{i})+\log m_{i}$ (1)
@@
  • $\log m_{i}$ is a correction factor accounting for the multiplicity $m_{i}$ of how many times the block $x_{i}$ appears.
For a generalised version of BDM holding for any encodable object, see [68].
The Coding Theorem Method (CTM) is a method based on the Coding Theorem and Algorithmic Probability [28, 58], which connects classical probability to algorithmic complexity [62, 64, 63]. The CTM maps sets of micro programs (e.g., small Turing machines) to small assembly objects for which it can empirically estimate the algorithmic probability $P(\cdot)$ of an object, such as a time series, based on the following relationship [69].
$-\log P(s)=K(s)=-\log\left(\mathbf{m}\left(s\right)\right)\pm\mathbf{O}(1)=-\log\left(\sum\limits_{p\in\left\{w\colon U(w)=s\right\}}2^{-|p|}\right)\pm\mathbf{O}(1)\text{ ,}$ (2)
@@
  • $P(s)$ is the algorithmic probability of string $s$;
  • $K(s)$ is the (prefix) algorithmic complexity of string $s$;
  • $\mathbf{m}\left(s\right)$ is a maximal semicomputable semimeasure on the object $s$;
  • $\sum\limits_{p\in\left\{w\colon U(w)=s\right\}}2^{-|p|}$ is the universal (a priori) probability of the event $s$.
Notice that a semicomputable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if for any other semicomputable semimeasure $\mu\left(\cdot\right)$—including any computable probability measure one may arbitrarily choose)—, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)$. The universal probability of an event can be understood as the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that generates the event.
CTM produces and stores the set of Gödel numbers that correspond to all the programs that compute an object, such as an integer sequence, up to the given digit or any other recursively describable [62, 64]. Each program can then be uncompressed from its unique (Gödel) number and run to produce the next digit for predictive purposes with the programs themselves the abstract future-planning models. While CTM operates by brute force, BDM leverages the pre-computed distributions that can be queried in linear time and stiches together longer explanations from small computer programs according to the rules of information theory to guide the search of the best sequence of programs explaining larger objects. In this sense, BDM can be thought of as a quintessential type of neural network transformers (as in self attention) where it estimates the local (short-range) causality through algorithmic complexity while computing long-range correlations through Shannon Entropy guaranteed convergence (worse case) [68].
On the one hand, CTM provides an approximation to algorithmic probability $P(s)$ by connecting the empirical frequency of occurrence of an object produced by a random computer program with its algorithmic complexity $K(s)$ and also keeps track of the set of programs that generated the original object, hence identifying the mechanistic generators.
On the other hand, BDM offers a method to map the micro programs produced by CTM to their corresponding pieces from the larger object to explain by decomposing the original object into smaller blocks for which micro programs have been found by CTM with a correction factor for block interactions (e.g. repetitions).
BDM allows for massive parallelisation. Objects with low complexity (i.e., higher causal impact at the global level) are the most frequent according to algorithmic probability and therefore are exponentially more frequent counteracting its intractability. BDM and CTM can be applied to test both:
@@
  • Prediction as planning: Using the BDM complexity as a proxy for the time series’ regularity, one can infer the predictability of future values. Lower BDM complexity implies a simpler underlying structure, which can help in forecasting future elements of the series—which is similar to how algorithmic probability and Levin’s universal distribution can be used for predictive modelling. (See Sections 10.3 and 10.4). This is related to planning, because once several program pathways are identified, one can verify each against the next token and update the program set (by discarding those programs that did not fit the next token) while keeping the shortest program criterion.
@@ 4.2.1 Why CTM and BDM as standard for abstraction and planning
  • CTM helps identify the set of candidate underlying generative mechanisms and provide a set of models from which it can actively predict future values by running it further into the future providing a set of projections. CTM forecasting requires an iterative refinement process in which multiple possible generative programs are tested and updated. CTM can help select the most likely program candidates from CTM by favouring those with lower complexity in accordance with the principles of algorithmic probability.
@@ 4.3 Comprehension via Algorithmic Probability
It should also be noticed that CTM/BDM is not purely a brute-force approach, requires no previous data and its current implementation required orders of magnitude less computational power. While CTM alone would be a brute-force approach that seeks the shortest computer programs explaining the data, BDM combines it with traditional pattern matching, meaning that CTM/BDM combines the best of both worlds, right at the fine balance between what traditional Machine Learning and Deep Learning approaches implement while also combining it with optimal Bayesian causal inference [35, 61]. We have called this approach Algorithmic Information Dynamics [71, 58, 59].
@@ 5.1 Next-digit Prediction Task with Binary and Non-binary Sequences
Figure 3 shows an increase in complexity as was expected, given the design of each group of generated sequences. The plot suggests that BDM can capture (and can generate) better complexity and randomness, since its values increase more consistently as complexity increases, unlike other measures. Shannon-entropy-based measures (and cognates) can account for statistical randomness only. Compression algorithms, for example, decrease as complexity increases, becoming more difficult to find regularities and increasing compression length as a function of complexity growth.
@@ 5.3 Code Generation Task with Non-binary Sequences
According to the results ( 11 top), the vast majority of correct cases are print failing to compress the sequences. This indicates that in most instances where the system correctly identifies a sequence, it does so by simply outputting the sequence as is, without any attempt at compression.
@@
Figure 12 (bottom) shows how no-compression generally increases with complexity, except for Mathematica, where the no-compression percentage is lower at complexity level 2 than at level 1. This happened because Mathematica has the capacity to computationally replicate several well-studied and known sequences of numbers. This capacity leads to shorter code at complexity level 2. However, at complexity level 3, the trend aligns with other languages, showing direct proportionality between complexity and no-compression.
@@ 6 SuperARC-seq
Based on the previous experiments, it is possible to characterise one test directly related to the SuperARC framework: the SuperARC-seq. The objective of this test is to quantify intelligence and related cognitive capacities, specifically, reasoning and comprehension, drawing inspiration from the work in [74] and the theoretical and empirical studies here introduced. As mentioned, this test is grounded in one of the fundamental cognitive tasks: recognising patterns and evaluating the complexity of finite sequences, which inherently requires a level of understanding in order to provide a meaningful explanation. In our experiment, we generated short integer sequences (100 binary and 90 integer-valued in general, as seen in subsections 10.15 and 10.16, respectively, in Suppl. Information) and tasked several advanced LLMs with deriving a formula capable of reproducing the each of the target sequences.
@@
    Ordinal: The model provided a mapping based on the indices where “1”s occur in the sequence. This response reflects an attempt by the model to analyse and map some logical structure to the sequence, making it more valuable than simply reproducing it verbatim. For integer sequences in general, a simple ASCII mapping was performed to convert from integers to binary encodings.
@@
For any given tested model, the percentages of results belonging to each group can be combined as a vector of results, $\rho=[\%_{c,np,no},\%_{c,o},\%_{c,p},\%_{inc}]$, such that $\sum\rho_{i}=1$ as the percentages will be represented in the range [0,1] to resemble probabilities. We know, beforehand, that the best performing model would be one with $\rho_{best}=[1,0,0,0]$. Thus, a first possible test would be to check the overall percentage of correct answers.
$\varphi_{a}=\sum_{i=1}^{3}\rho_{i},$ (3)
which would range from 0 to 1 for models that are not able to reproduce any sequence to models which perfectly reconstruct the sequences, respectively. However, this only accounts for the ability of LLMs to reproduce the initial sequence (planning) but not for their compression capabilities. To account for the latter, let us assume that the best possible algorithm for each element of the data set is $\mathcal{B}_{k,j}$, such that $\mathcal{B}_{k,j}()=D_{k,encoded}[j]$, and here the algorithm does not have a particular input, similar to the definition of algorithmic complexity. Thus:
@@
due to the information non-increase theorem and to the fact that no inputs were used in the function. The ratio $K(D_{k,encoded}[j])/K(\mathcal{B}_{k,j})$ consistently falls within the range [0,1] for medium to long sequences when no embedding algorithms are employed. This behavior arises because approximations of algorithmic complexity are less reliable for short sequences, primarily due to the overhead inherent in theoretical computations. In order to surpass this limitation, since the difference between the true algorithmic complexity value and its approximation is bounded by a linear constant in general, instead of assessing the absolute algorithmic complexity (or any of its approximations), we shall consider a normalized version of it. To approximate algorithmic complexity, we will use the BDM/CTM approach, as described in detail in previous sections.
To build a normalized version of BDM/CTM, as pointed out in previous works [70], for any object of arbitrary size, it is possible to construct analogous objects that attain the minimum and maximum possible values of algorithmic complexity according to the Block Decomposition Method (BDM):
  • Minimum complexity object: This case is straightforward and corresponds to an object composed entirely of a single repeating symbol - for instance, a binary string consisting solely of zeros.
  • Maximum complexity object. The maximum BDM value is achieved by an object whose decomposition (according to a specified algorithm) results in slices that exhibit the highest values of the Coding Theorem Method (CTM), with each distinct slice occurring only once until all possible configurations of the given shape have been exhausted.
The primary advantage of considering a normalized measure lies in its ability to enable comparisons between objects of varying sizes, effectively mitigating the influence of size on the measure itself. This property is particularly in the case of the present study, where we compare complexities of sequences and formulas generating them.
This way, the following ratio presents itself as an interesting weighting factor for the probabilities in equation 3:
$nBDM(D_{k,encoded}[j])/nBDM(\mathcal{B}_{k,j})$ (5)
The ratio in equation 5 measures how the algorithmic complexity of the formula and sequence compare to the other possible outputs of the LLM. If relative algorithmic complexity (measured by the normalized BDM value) for the formula is greater than it was for the sequence itself, this suggests the LLM did not success in compressing the input sequence (it made the formula have a greater relative algorithmic complexity). On the other hand, if the opposite occurs, then the LLM could compress the sequence comparatively to other possible outputs of the LLM. The ratio in equation 5 ranges from 0 to a positive value $M>1$, which happens when the best possible compression is achieved (the inverse mapping of CTM). Since $M$ is not known beforehand, we can use a nonlinear mapping that saturates the value of the ratio to a maximum value of 1 (similar to an activation function). The hyperbolic tangent function can be used in this case, since $\text{tahn}(0)=0$ and $\lim_{x\to\infty}\text{tahn}(x)=1$. Thus, a candidate weighting factor for the probabilities in 3 is:
$\delta_{k,j}=\text{tanh}\left(\frac{nBDM(D_{k,encoded}[j])}{nBDM(\mathcal{B}_{k,j})}\right)$ (6)
with the best possible value of $\delta_{k,j}$ approaching 1 in a perfect compression scenario. Since we have several algorithms classified under each of the four types (according to their structure), instead of using the individual ratios for each type $k$, we shall use the harmonic mean per type, defined as:
$\delta_{k}=\frac{n_{k}}{\displaystyle\sum_{j=1}^{n_{k}}\delta_{k,j}}\text{ for}R_{k,j}\text{ of type }k,$ (7)
@@
$\varphi_{b}=\sum_{i=1}^{3}\delta_{i}\rho_{i}.$ (8)
Deliberately, we want to privilege models that do not simply copy or provide ordinal mappings of the input sequences. Thus we can attribute higher weights to types that are correct and do not copy nor print the results. We also want to give more weight to programs that provide ordinal mappings when compared to print cases. Then, considering a power-law weighting strategy, the final test metric is:
$\varphi=\delta_{1}\rho_{1}+\frac{\delta_{2}\rho_{2}}{10}+\frac{\delta_{3}\rho_{3}}{100}.$ (9)
It can be seen that $\varphi\in[0,1]$ encompasses different behaviours. For example, $\varphi\in[0,0.01]$ if only print-type models are outputted. Also, $\varphi\in[0,0.1]$ if only ordinal-like formulas are created. Finally, $\varphi\in[0,1]$ in cases where the LLMs create formulas that are always correct, do not copy nor create ordinal mappings. The ranges will be populated with varying compression levels corresponding to the algorithms obtained. Overall, if the score is 0, all the formulas were wrong. If it is 0.5, it can represent the case where half the outputs were correct and half wrong, with the formulas produced with highest compression levels. So, in a regular half and half case, since compression will not be optimal, the test score is less than 0.5.The test performance results for each model are calculated using equation 9 for $\mathcal{T}$ in Algorithm 1.
@@
As described, equation 9 tests for two features, compression via non-print computer programs and non-ordinal mathematical formulas to the input sequence, and prediction, by running all programs and all formulas to match each sequence digit, and penalising them when they did not represent an actual compressed model that generated a possible new digit of the sequence when run in reverse, i.e. when ‘decompressed’. The test formula assigns greater importance to correct cases that are not solutions of the type ‘print($s$)’ where $s$ is the sequence for which the AI system is asked for a model, given that a print model does not allow generalisation by prediction through simulation, as running a print command will only print up to the last digit. The same is true for what we call ‘ordinals’, which is simply indicating the index of the non-zero non-one element in the binary (or binary embedded) sequence, meaning that, together with the ‘print’ case, the system failed in its attempts at abstracting features of the object. Finally, the formula punishes ordinal and print answers in a weighted fashion. The best performer can only reach a $\varphi$ of 1 while the lowest value is 0.